A Complete Bibliography of Scientific American: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 May 2018 Version 1.05

Title word cross-reference

g [?].

3000 [?].

600-Foot [?].

Aberration [?]. Aborigine [?]. Abortion [?]. Absolute [?, ?]. Acceleration [?]. Accelerator [?, ?]. Accelerators [?]. Acetabularia [?]. Acid [?, ?]. Acids [?, ?]. Acoustic [?]. Acoustical [?]. Acoustics [?]. ACTH [?]. Action [?]. Activation [?]. Acute [?]. Adaptations [?]. Add [?]. Addiction [?]. Adhesives [?]. Adjustable [?]. Advances [?, ?, ?, ?]. Aerial [?]. Aerodynamics [?]. Affair [?]. Africa [?, ?]. Aftereffects [?]. Afterimages [?]. Age [?, ?, ?, ?]. Aging [?, ?, ?, ?]. Ago [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Ago [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Agricultural [?]. Agriculture [?]. Air [?, ?, ?, ?, ?]. Air-Breathing [?]. Air-Conditioned [?]. Air-Traffic [?]. Airborne [?]. Aircraft [?, ?]. Airglow [?]. Alchemy [?]. Alessandro [?]. Algae [?]. Algebra [?]. All-Weather [?]. Alps [?]. Altitude [?]. Amateur [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Amateur [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. America [?, ?, ?]. American [?]. Amoebae [?]. Amoeboid [?]. Amorphous [?, ?]. Amorphous-Semiconductor [?, ?]. Amphibian [?]. Analgesic [?]. Analysis [?, ?, ?, ?, ?, ?]. Analytic [?]. Analyzed [?]. Ancient [?, ?, ?, ?, ?, ?, ?]. Andes [?]. Animal [?, ?, ?]. Animals [?, ?, ?, ?, ?, ?, ?, ?]. Annual [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Antarctic [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Anti [?]. Anti-Ballistic-Missile [?]. Antibiotics [?]. Antibodies [?, ?]. Antigens [?]. Antimatter [?]. Antiquity [?]. Antlers [?]. Anxiety [?, ?]. Apes [?]. Apparatus [?]. Applications [?, ?, ?]. Approaching [?]. Aquatic [?]. Arabia [?]. Ararat [?]. Archer [?]. Architectural [?, ?]. Architecture [?]. Arctic [?]. Arithmetic [?]. Armaments [?]. Arms [?]. Arousal [?]. Arrival [?]. Art [?, ?]. Artemis [?]. Artificial [?, ?, ?, ?]. Asparagine [?]. Asphyxia [?]. Aspirations [?]. Aspirin [?]. Assassins [?]. Associate [?]. Assyrian [?]. Astroblemes [?]. Astronomy [?, ?, ?, ?, ?, ?, ?]. Astrophysics [?]. Atherosclerosis [?]. Atmosphere [?, ?, ?, ?, ?, ?]. Atmospheres [?]. Atmospheric [?]. Atom [?]. Atomic [?]. Attack [?]. Attention [?]. Attitude [?]. Attitudes [?]. Attractants [?]. Auditory [?]. Aurora [?]. Australian [?]. Authors [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Authors [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Autobiographies [?]. Automatic [?]. Automobile [?, ?]. Avalanches [?]. Axon [?].

B.C. [?]. Babies [?]. Baboons [?]. Backward [?]. Bacterial [?, ?, ?, ?, ?, ?]. Ball [?]. Ballistic [?]. Balloon [?]. Ban [?]. Battle [?]. Beaches [?]. Beams [?]. Bearings [?]. Bears [?]. Becomes [?]. Bee [?]. Bees [?]. before [?]. Beginnings [?]. Behavior [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Behavioral [?, ?]. Belts [?]. Beneficiation [?]. Bering [?]. between [?, ?]. Beyond [?]. Bibliography [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Bibliography [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Bilingualism [?]. Biochemical [?]. Biochemistry [?, ?]. Biological [?, ?, ?]. Biology [?]. Biotin [?]. Bird [?]. Birds [?, ?, ?]. Birth [?]. Bison [?]. Black [?]. Blindness [?]. Blood [?, ?]. Blood-Vessel [?]. Blue [?]. Blue-Green [?]. Body [?]. Bohr [?, ?]. Bone [?]. Book [?]. Books [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Books [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Boom [?]. Boomerangs [?]. Boron [?, ?]. Bowerbirds [?]. Boyle [?]. Brahe [?]. Brain [?, ?, ?, ?, ?, ?, ?, ?]. Brauron [?]. Brazil [?]. Breath [?]. Breathing [?, ?, ?]. Breeder [?, ?]. Bridge [?]. Brownian [?]. Buffalo [?]. Building [?, ?]. Buoyancy [?]. Butterflies [?].

Calcutta [?]. Cambrian [?, ?]. Camp [?]. Campsite [?]. Can [?, ?]. Cancer [?, ?]. Cantorian [?]. Capital [?, ?]. Carbon [?, ?]. Care [?, ?]. Cargo [?]. Cargo-Handling [?]. Casting [?]. Cataracts [?]. Cave [?]. Celestial [?]. Cell [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Cells [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Cellular [?]. Cellulose [?]. Census [?, ?]. Centered [?]. Central [?]. Ceramics [?]. Cerebral [?]. Chamber [?, ?]. Chance [?]. Change [?, ?, ?]. Changing [?, ?]. Channeling [?]. Chelation [?]. Chemical [?, ?, ?, ?, ?, ?, ?, ?, ?]. Chemicals [?]. Chemistry [?, ?, ?, ?, ?, ?, ?, ?, ?]. Chimpanzees [?, ?]. China [?]. Chinampas [?]. Chinese [?]. Chingis [?]. Chlorophyll [?]. Chondrites [?]. Chondrules [?]. Chromatography [?]. Chromosome [?, ?, ?]. Chromosomes [?, ?]. Chromosphere [?]. Cilia [?]. Circulation [?, ?]. Cities [?, ?, ?, ?, ?, ?, ?]. Citizen [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Citizen [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. City [?, ?, ?, ?]. Ciudad [?]. Civilization [?, ?]. Clay [?]. Cleaning [?]. Cliff [?]. Climate [?, ?]. Clock [?]. Clocks [?]. Clotting [?]. Cloud [?]. Clouds [?, ?]. Coanda [?]. Code [?, ?, ?, ?, ?, ?]. Codes [?]. Cognitive [?]. Cold [?, ?]. Collagen [?, ?]. Collapse [?]. Color [?, ?, ?]. Colors [?]. Columbian [?]. Combustion [?]. Commissure [?]. Common [?]. Communicate [?, ?]. Communication [?, ?, ?]. Company [?]. Competition [?]. Composite [?]. Composition [?]. Compounds [?]. Computer [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Computers [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Concepts [?]. Conceptual [?]. Concrete [?].

Conditioned [?]. Conduction [?]. Confirmation [?]. Conflict [?]. Conformity [?]. Conquests [?]. Conservation [?]. Constructed [?]. Contact [?, ?]. Continental [?, ?, ?, ?]. Continuous [?]. Contraction [?]. Control [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Controlled [?, ?]. Copernican [?]. Copper [?]. Corals [?]. Corona [?]. Coronary [?]. Correcting [?]. Corrosion [?]. Corrugate [?]. Cortex [?]. Cosmic [?, ?, ?, ?]. Cosmology [?]. Cost [?]. Count [?]. Criminal [?]. Crises [?]. Cromwell [?]. Cryptology [?]. Crystal [?]. crystallography [?]. Crystals [?, ?, ?, ?, ?, ?, ?]. Cult [?]. Cultivation [?]. Culture [?, ?]. Cumae [?]. Current [?]. Currents [?]. curve [?]. Cycle [?]. Cycles [?].

Daddy [?]. Damage [?]. Dance [?]. Data [?]. Davy [?]. Day [?]. Death [?, ?]. Decline [?]. Deep [?, ?]. Deep-Ocean [?]. Defective [?]. Delayed [?]. Demon [?]. Density [?, ?]. Deprivation [?]. Depth [?]. Dermatoglyphics [?]. Desalting [?]. Desegregation [?]. Desert [?, ?, ?]. Detected [?]. Detection [?, ?]. Detector [?]. Detectors [?]. Development [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Devices [?]. Diagnosis [?]. Dialects [?]. Diamonds [?]. Diesel [?]. Differences [?, ?]. Differentiation [?, ?]. Digesting [?]. Digits [?]. Dimensional [?, ?]. Dimensions [?]. Dioxide [?]. Disadvantaged [?]. Disarmament [?, ?]. Discovery [?, ?, ?]. Disease [?, ?, ?, ?, ?]. Diseases [?]. Dislocations [?]. Dissonance [?]. Distances [?]. Distribution [?, ?]. Divide [?]. Diving [?]. DNA [?, ?, ?, ?]. do [?]. Doves [?]. Dreaming [?]. Drift [?, ?, ?, ?]. Drug [?]. Drugs [?, ?, ?]. Ductile [?]. Dust [?]. Dwarf [?]. Dwellers [?]. Dynamics [?, ?]. Dynamo [?].

Earlier [?]. Earliest [?]. Early [?, ?, ?, ?, ?, ?, ?, ?, ?]. Earth [?, ?, ?, ?, ?, ?]. East [?]. Echoes [?]. Ecological [?, ?, ?]. Ecology [?, ?]. Economic [?, ?, ?]. Economics [?, ?]. Economy [?]. edited [?]. Education [?, ?]. Effect [?, ?]. Effects [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Eggs [?]. Eidetic [?]. Eland [?]. Electric [?, ?, ?, ?, ?]. Electrical [?, ?, ?]. Electrically [?]. Electricity [?]. Electrocardiogram [?]. Electromagnetic [?]. Electron [?, ?, ?, ?]. Elements [?, ?, ?]. Elephant [?]. Elephant-Hunting [?]. ellipse [?]. Embryological [?]. Emission [?]. Emitting [?, ?]. Emotional [?]. Empire [?]. End [?]. Endotoxins [?]. Enduring [?]. Energetics [?, ?]. Energy [?, ?, ?, ?]. Engine [?, ?, ?]. Engines [?]. Entry [?]. Environment [?, ?]. Enzyme [?]. Enzymes [?, ?]. Eötvös [?]. Eradication [?]. Erectus [?]. Erratum [?, ?, ?, ?]. Error [?]. Error-Correcting [?]. Escape [?]. Esteem [?]. Etruscans [?]. Euclid [?]. Euclidean [?]. Eupalinus [?]. Eutectics [?]. Evolution [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. examination [?]. Excavation [?]. Excavations [?]. Exercise [?]. Expectations [?]. Experience [?]. Experiment [?, ?, ?, ?]. Experimental [?, ?]. Experiments [?, ?, ?, ?, ?, ?, ?, ?]. Exploding [?, ?, ?]. Exploration [?, ?]. Explorer [?].

Explosions [?, ?]. Expressions [?]. Eye [?, ?, ?].

Facial [?]. Factor [?]. Factors [?]. Failure [?, ?]. Fall [?]. False [?]. Faraday [?]. Fast [?, ?]. Fast-Neutron [?]. Faster [?]. Fat [?, ?]. Feather [?]. Feel [?]. Felix [?]. Fermi [?]. Ferrites [?]. Fertility [?]. Fertilizers [?]. Fever [?]. Fiber [?, ?]. Fiber-Reinforced [?]. Fibrinogen [?]. Field [?, ?, ?]. Fields [?, ?, ?, ?, ?]. Fighting [?]. Films [?]. Fine [?]. Fire [?]. Fireball [?]. First [?, ?, ?]. Fish [?, ?]. Fishes [?, ?, ?, ?]. Fission [?]. Fixed [?]. Fixed-Point [?]. Flare [?]. Flash [?, ?]. Fleas [?]. Fleming [?]. Flies [?]. Flight [?, ?, ?, ?]. Flight-Control [?]. Floaters [?]. Floor [?, ?, ?, ?, ?, ?]. Flow [?]. Flower [?]. Fluid [?, ?]. Fluidization [?]. Fluids [?]. Fly [?, ?]. Fog [?]. Food [?, ?, ?]. Foot [?]. Footprints [?]. Force [?, ?]. Foreign [?]. Forest [?]. Forests [?]. Forgetting [?, ?]. Forgotten [?]. Form [?]. Formed [?]. Fossil [?]. Fossils [?]. Foundations [?]. Foxglove [?]. Fractionating [?]. Fracture [?]. Fragments [?]. Free [?]. Freezing [?]. Friendly [?]. Frog [?]. Frogs [?]. Frozen [?]. Fruit [?]. Fungi [?]. Fungus [?]. Fusion [?, ?].

Galaxies [?, ?, ?, ?, ?, ?, ?]. Galaxy [?]. Galvanomagnetic [?]. Game [?]. Games [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Games [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Gamma [?]. Gardens [?]. Gas [?, ?]. Gases [?, ?]. Gateway [?]. Gene [?, ?]. General [?]. Generation [?, ?]. Genes [?, ?, ?]. Genetic [?, ?, ?, ?, ?, ?, ?, ?, ?]. Genetics [?]. Geological [?]. Geology [?, ?]. Geomagnetic [?]. Geometry [?, ?]. George [?]. Germ [?]. Germ-Free [?]. German [?]. get [?]. Ghetto [?]. Giant [?]. Gland [?, ?, ?, ?]. Glass [?, ?]. Glasses [?]. go [?, ?]. Goggles [?]. Golgi [?]. Gradients [?]. Grains [?]. Gravitational [?]. Gravity [?]. Great [?, ?, ?, ?, ?]. Greece [?]. Greeks [?]. Green [?, ?, ?]. Ground [?]. Growth [?, ?, ?, ?, ?, ?]. Guayana [?]. Gulf [?]. Gulls [?, ?].

Habitat [?]. Hacilar [?]. Hagfish [?]. Hallucinogenic [?]. Hamster [?]. Hand [?]. Handling [?]. Harappans [?]. Harvesting [?]. Healing [?]. Health [?, ?]. Heart [?, ?, ?, ?, ?]. Heat [?, ?, ?, ?]. Heavy [?]. Hebrews [?]. Helicopter [?]. Helium [?, ?]. Hemoglobin [?, ?]. Hemophilia [?]. Herding [?]. Hibernators [?]. Hidden [?]. High [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. High-Energy [?]. High-Power [?]. High-Pressure [?]. High-Speed [?, ?]. High-Temperature [?]. High-Voltage [?]. Hill [?]. Himalayas [?]. History [?]. Holography [?, ?]. Homo [?, ?]. Honeybees [?]. Hopewell [?]. Hormone [?, ?]. Hormones [?, ?, ?]. Horns [?]. House [?]. Human [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Humphry [?]. Hunter [?]. Hunting [?]. Hurricane [?]. Husbandry [?]. Hybrid [?, ?, ?]. Hydrogen [?]. Hydroxyl [?, ?]. Hypernuclei [?]. Hypersensitivity [?].

Icarus [?]. Ice [?, ?, ?, ?]. II [?, ?, ?]. III [?, ?, ?]. Illusion [?, ?]. Illusions [?]. Images [?, ?, ?, ?, ?]. Imitating [?]. Immunity [?]. Immunoelectrophoresis [?]. Impact [?, ?]. Inclusion [?]. Index [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. India [?, ?, ?]. Indian [?, ?]. Indies [?]. Induction [?]. Industrial [?]. Infancy [?]. Infants [?]. Infections [?]. Infectious [?]. Influence [?]. Information [?, ?, ?]. Infra [?]. Infra-Cambrian [?]. Infrared [?, ?, ?]. Inhibition [?, ?]. Inputs [?]. Insect [?, ?]. Insects [?, ?, ?, ?]. Instinct [?]. Instruments [?]. Integrated [?]. Intelligence [?, ?]. Intense [?]. Intensive [?]. Interacting [?]. Interaction [?]. Interacts [?, ?]. Interference [?]. Interferon [?]. Interior [?]. Interplanetary [?]. Interstellar [?]. Iran [?]. Iron [?]. Ishango [?]. Isimila [?]. Isolation [?]. Isolators [?]. Isomers [?]. Italian [?]. IV [?, ?, ?, ?].

Jammer [?]. Japan [?]. Jerusalem [?]. Jet [?]. Jordan [?]. Josephson [?]. Judgments [?]. jumps [?]. Jupiter [?, ?].

Keratins [?]. Khan [?]. Kidney [?]. Kill [?]. Kilomegacycle [?]. King [?]. Kinins [?]. Kinship [?]. Korea [?].

L [?]. L-Asparagine [?]. Laboratory [?, ?]. Lacrimal [?]. Lake [?]. Lakes [?]. Land [?, ?, ?]. Landing [?]. Language [?, ?]. Large [?]. Laser [?, ?, ?, ?, ?, ?]. Lasers [?, ?, ?, ?]. Last [?]. Lateritic [?]. Lathe [?]. Law [?]. Laws [?]. Layer [?]. Layers [?]. Leakage [?]. Learned [?]. Learning [?, ?]. Lee [?]. Leibniz [?]. Length [?]. Lesson [?]. Letters [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Letters [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Leukemia [?, ?]. Level [?]. Lichens [?]. lies [?]. Life [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Light [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Light-Emitting [?]. Lightning [?]. Lily [?]. Lines [?]. Liquid [?, ?, ?, ?]. Liquids [?, ?]. Lithium [?]. Lives [?]. Living [?, ?, ?, ?, ?, ?, ?]. Localization [?]. Locating [?]. Location [?]. Locust [?]. Logic [?, ?]. Loglan [?]. Longest [?]. Longlegs [?]. Lovebirds [?]. Low [?]. Low-Altitude [?]. Luminescence [?, ?]. Luminous [?]. Lunar [?]. Lung [?]. Lungs [?]. Lymphatic [?]. Lysenko [?]. Lysosome [?]. Lysosomes [?]. Lysozyme [?].

Macedonia [?]. Machine [?, ?]. Machines [?, ?]. Made [?]. Magellan [?]. Magnetic [?, ?, ?, ?, ?, ?, ?, ?, ?]. Magnetism [?, ?]. Magnetometer [?]. Magnetosphere [?]. Magnetothermoelectricity [?]. Magnets [?, ?]. Make [?, ?]. Making [?]. Malaria [?]. Mammalian [?]. Mammals [?]. Mammoth [?]. Man [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Man-Made [?]. Man-of-War [?]. Manipulators [?]. Mantle [?]. Marihuana [?]. Marine [?]. Mariner [?, ?, ?, ?]. Markets [?]. Marking [?]. Mars [?]. Masers [?, ?]. Master [?]. Materials [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mathematical [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mathematical [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mathematician [?]. Mathematics [?, ?, ?, ?, ?]. Matter [?, ?, ?, ?]. Maturation [?]. Max [?]. Maxwell [?, ?]. McGraw [?]. McGraw-Hill [?]. McKinley [?]. Meanders [?]. Measles [?]. Measurement [?, ?, ?]. Measuring [?]. Mechanical [?]. mechanics [?, ?]. Mechanism [?, ?, ?, ?]. Mechanisms [?]. Medical [?, ?, ?, ?]. Medicine [?]. Mediterranean [?]. Meeting [?]. Mekong [?]. Membrane [?, ?, ?]. Membranes [?]. Memories [?]. Memory [?, ?, ?]. Messenger [?]. Metabolism [?]. Metal [?]. Metalliding [?]. Metallurgy [?]. Metals [?, ?, ?, ?, ?]. Metamorphosis [?]. Meteorites [?]. Meteoritic [?]. Methods [?, ?, ?, ?]. Metropolis [?, ?]. Metropolitan [?, ?]. Mexico [?]. Michelson [?]. Microelectronics [?]. Micrometeorology [?]. Micropaleontology [?]. Microwaves [?]. Midgets [?]. Migration [?, ?]. Mile [?]. Military [?]. Milk [?]. Mind [?]. Minerals [?, ?]. Minus [?]. Missile [?]. Missions [?, ?]. Misuse [?]. Mitochondrion [?, ?]. Model [?]. Model-Building [?]. Modern [?, ?, ?]. Modification [?]. Modulation [?, ?]. Moiré [?]. Molds [?]. Molecular [?, ?, ?]. Molecule [?, ?, ?, ?]. Molecules [?, ?]. Money [?]. Mongol [?]. Monkey [?]. Monkeys [?, ?]. Monomolecular [?]. Monopoles [?]. Monstrosus [?]. Moon [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Moore [?]. Morley [?]. Mosaics [?]. Mosquito [?]. Mössbauer [?]. Moths [?]. Motion [?, ?]. Motivation [?]. Motor [?]. Mouse [?]. Move [?]. Movement [?, ?]. Movements [?]. Moves [?]. Muon [?]. Muonium [?]. Muscle [?]. Muscles [?]. Muscular [?].

Nabataeans [?]. Narcotic [?]. National [?, ?]. Nationality [?]. Natural [?, ?]. Nature [?, ?, ?, ?, ?, ?, ?, ?]. Navigation [?, ?, ?]. Nebula [?]. Nebulae [?]. Needs [?]. Negro [?]. Neolithic [?, ?, ?, ?]. Nerve [?, ?]. Nervous [?, ?]. Nests [?]. Neurophysiology [?]. Neutral [?]. Neutrino [?, ?]. Neutrinos [?, ?]. Neutron [?, ?, ?]. Newer [?]. Nice [?]. Niels [?, ?]. Nigeria [?]. Night [?]. Ninian’s [?]. Nobel [?]. Noble [?, ?]. Noctilucent [?]. Noise [?, ?]. Non [?, ?]. Non-Cantorian [?]. Non-Euclidean [?]. Nonuniform [?]. North [?]. Nuclear [?, ?, ?, ?, ?, ?]. Nuclear-Test [?]. Nuclei [?, ?]. Nucleic [?, ?, ?, ?]. Nucleotide [?].

Number [?]. Numerical [?]. Numismatics [?].

Objects [?]. Observations [?]. Observatories [?]. Observing [?, ?]. Obsidian [?]. Ocean [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Oceanic [?, ?, ?]. Oceans [?]. Octopus [?]. Odor [?]. Oil [?]. Olympic [?]. Omega [?]. Omega-Minus [?]. Open [?]. Open-Heart [?]. Operation [?]. Opiates [?]. Optical [?, ?, ?, ?]. Optics [?]. Orbiter [?]. Orchids [?]. Ordinary [?]. Ores [?, ?]. Organic [?, ?, ?]. Organizations [?]. Origin [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Origins [?, ?, ?, ?, ?, ?, ?]. Orion [?]. Orthodox [?]. Oryx [?]. Other [?]. Outpost [?]. Outputs [?]. Oxygen [?].

Pacemaker [?]. Pacific [?]. Pain [?]. Palace [?]. Paleo [?]. Paleo-Indian [?]. Paleolithic [?, ?, ?]. Paleontological [?]. Paradox [?, ?, ?]. Parathyroid [?]. Park [?]. Particle [?, ?]. Particle-Detectors [?]. Particles [?, ?, ?, ?]. Patent [?]. Path [?]. Pathology [?]. Pattern [?]. Patterns [?, ?]. Peasant [?]. Peculiar [?, ?]. Pella [?]. Penguins [?]. Penicillins [?]. People [?]. Perception [?, ?, ?, ?, ?, ?, ?]. Perpetual [?]. Persian [?]. Peru [?]. Pesticides [?]. Petrified [?]. Petroglyphs [?]. Petroleum [?, ?]. Phalarope [?]. Phase [?]. Phases [?]. Pheromones [?]. Photograph [?]. Photographs [?]. Photography [?]. Photolysis [?]. Photon [?]. Photosynthesis [?, ?, ?, ?]. Physical [?, ?]. Physicist [?]. Physics [?, ?, ?, ?, ?]. Physiological [?]. Physiology [?, ?, ?]. Piano [?]. Picture [?]. Pigment [?]. Pigments [?, ?]. Pigs [?]. Pineal [?]. Pinocytosis [?]. Pipe [?]. Pipelines [?]. Place [?]. Place-Learning [?]. Plague [?]. Plan [?, ?]. Planetary [?]. Planets [?, ?]. Planned [?, ?]. Planning [?]. Plant [?, ?, ?]. Plants [?, ?, ?, ?]. Plasma [?]. Plasmas [?]. Plastic [?]. Plasticity [?]. Plastics [?]. Platelets [?]. Pleiades [?]. Point [?]. Polar [?]. Polarized [?]. Policy [?]. Polishing [?]. Pollen [?]. Pollutants [?]. Pollution [?, ?, ?]. Polyethylene [?]. Polygraph [?]. Polymeric [?]. Polymers [?]. Polyoma [?]. Polyribosomes [?]. Poor [?]. Population [?, ?, ?, ?, ?, ?]. Pores [?]. Porphyria [?]. Portuguese [?]. Possessions [?]. Potential [?]. Poultry [?]. Poverty [?, ?, ?]. Power [?, ?, ?, ?, ?]. Pre [?, ?]. Pre-Cambrian [?]. Pre-Columbian [?]. Precisely [?]. Predatory [?]. Prehistoric [?, ?, ?]. Prehistory [?, ?]. Premature [?]. Present [?]. President [?]. Pressure [?, ?]. Prevention [?]. Primeval [?]. Primitive [?]. Prizes [?]. Probability [?]. Problem [?, ?, ?]. Problem-Solving [?]. Process [?]. Processes [?]. Processing [?, ?]. Production [?, ?]. Programming [?]. Programs [?]. Progress [?]. Project [?]. Proof [?]. Properties [?, ?, ?, ?, ?]. Propulsion [?]. Protein [?, ?, ?, ?, ?, ?]. Protein-Digesting [?]. Proteins [?, ?, ?]. Protopsychology [?]. Psychiatry [?]. Psychological [?]. Psychophysics [?]. Psychotherapy [?]. Public [?]. Puffs [?]. Pulsars [?]. Pulse [?]. Pulse-Code [?]. Pumping [?]. Pumps [?]. Pupil [?]. Purple [?].

Pygmies [?]. Pythagoreans [?].

Qanats [?]. Quantized [?]. Quantum [?, ?]. Quasi [?, ?, ?]. Quasi-Particles [?]. Quasi-Stellar [?, ?]. Queues [?]. Quick [?]. Quiet [?].

Rabbits [?]. Race [?]. Radar [?, ?]. Radiation [?, ?, ?, ?, ?]. Radiation-Imitating [?]. Radicals [?, ?]. Radio [?, ?, ?, ?, ?, ?, ?]. Radio-Emitting [?]. Radiography [?]. Ranger [?]. Rapid [?]. Rational [?]. Rats [?]. Ray [?, ?, ?, ?]. Rays [?, ?, ?]. Re [?]. Re-Entry [?]. Reactions [?]. Reactors [?, ?, ?]. Reapportionment [?]. Receive [?]. Received [?]. Receptor [?]. Receptors [?]. Reclamation [?]. Recognition [?]. Recording [?]. Recovery [?]. rectangle [?]. Red [?, ?]. Red-Feather [?]. Redistricting [?]. Redundancy [?]. Refining [?]. Refrigeration [?]. Region [?]. Reinforced [?]. Relatives [?]. Relativism [?]. Relaxation [?]. Remembering [?]. Remote [?]. Renewal [?]. Repair [?]. Reproduction [?]. Reproductive [?]. Residential [?]. Resistance [?]. Resonance [?, ?]. Resonant [?, ?]. Resources [?, ?, ?]. Respiratory [?]. Reticulum [?]. Retina [?]. Retinal [?]. Retrieval [?]. Reversal [?, ?]. Reversals [?, ?]. Revolution [?, ?, ?]. Rhesus [?]. Rheumatic [?]. Ribosomes [?]. Ridged [?]. Ridges [?]. Rift [?]. Ring [?]. Rings [?, ?]. Rioters [?]. Rise [?, ?, ?]. River [?, ?]. Rivers [?]. RNA [?]. Roads [?]. Robert [?]. Rockets [?]. Role [?, ?]. Room [?]. Rotary [?]. Rotation [?, ?]. Royal [?]. Rozental [?]. Rudolf [?]. Rumford [?]. Run [?]. Ruth [?].

S [?]. Sahara [?]. Sailing [?]. Salamis [?]. Salmon [?]. Salpa [?]. Salt [?, ?]. Salt-Water [?]. Sanctuary [?]. Sand [?]. Sands [?]. Sap [?]. Sarcoplasmic [?]. Sardis [?]. Satellite [?]. Satellites [?, ?, ?]. Scale [?]. Scattering [?, ?]. Schizophrenia [?]. Schooling [?]. Science [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Science [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Sciences [?, ?, ?]. Scientific [?, ?, ?]. Scientist [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Scientist [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Scorpions [?]. Screw [?]. Screw-Worm [?]. Scythians [?]. Sea [?, ?, ?, ?, ?, ?, ?]. Sea-Floor [?, ?]. Seal [?]. Search [?]. Secondary [?]. Security [?, ?]. see [?]. Segregation [?]. Seismic [?]. Selection [?]. Self [?]. Self-Esteem [?]. Semiconductor [?, ?, ?]. Semiconductors [?]. Senses [?]. Sensing [?]. Sensory [?].

Sensory-Motor [?]. Sequence [?]. Set [?]. Settlements [?]. Sex [?, ?]. Sexual [?]. Seyfert [?]. Shadows [?]. Shales [?]. Shape [?, ?, ?]. Sharing [?]. Sharks [?]. Shelter [?]. Shelter-Centered [?]. Shelves [?]. Shock [?, ?]. Shooting [?]. Short [?]. Short-Term [?]. Siberia [?]. Signals [?]. Sing [?]. Single [?]. Single-Stranded [?]. Site [?, ?, ?]. Size [?, ?, ?]. Skin [?, ?, ?, ?]. Skua [?]. Sky [?]. Sleep [?]. Slime [?]. Small [?]. Smallest [?]. Smoking [?]. Snow [?, ?, ?]. Soaring [?]. Social [?, ?, ?, ?, ?, ?, ?, ?]. Society [?, ?]. Sociology [?]. Soil [?]. Soils [?]. Solar [?, ?, ?, ?, ?, ?]. Solid [?, ?, ?, ?, ?]. Solid-State [?]. Solids [?, ?, ?, ?, ?, ?]. Solutrean [?]. Solvated [?]. Solving [?]. Somatic [?]. Sonic [?]. Sound [?, ?]. Source [?]. Sources [?, ?]. South [?, ?]. Space [?, ?, ?, ?, ?, ?]. Span [?]. Spark [?]. Specialize [?]. Spectroscopies [?]. Spectroscopy [?]. Speech [?, ?]. Speed [?, ?, ?]. Spherules [?]. Spider [?]. Spills [?]. Split [?]. Spot [?]. Spreading [?, ?]. Squatter [?]. Squirrels [?]. St. [?]. Stabilized [?]. Standards [?]. Staphylococci [?]. Stapling [?]. Stars [?, ?, ?, ?, ?, ?, ?]. Start [?]. State [?, ?, ?]. States [?]. Steam [?, ?]. Steel [?, ?]. Steelmaking [?]. Steels [?]. Stellar [?, ?, ?, ?]. Steps [?]. Stereochemical [?]. Stimulation [?, ?]. Stimuli [?]. Stirling [?]. Stockholm [?]. Stone [?, ?]. Stones [?]. Storage [?, ?]. Straight [?]. Strains [?]. Strait [?]. Stranded [?]. Streamer [?]. Streams [?]. Strength [?]. Stress [?]. Stress-Corrosion [?]. String [?]. Strong [?]. Strongly [?]. Structure [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Studies [?]. Study [?, ?, ?, ?]. Subdwarf [?]. Subsidence [?]. Subsistence [?]. Substances [?]. Sun [?, ?, ?, ?]. Sunburn [?]. Superconducting [?, ?, ?]. Superconductivity [?, ?]. Superconductors [?]. superellipse [?]. Superfluid [?]. Superfluidity [?]. Superplastic [?]. Supersonic [?]. Support [?]. Surface [?, ?, ?, ?]. Surfaces [?]. Surgery [?, ?, ?]. Surgical [?]. Swimming [?]. Swiss [?]. Switch [?]. Switching [?, ?, ?]. Symbiosis [?]. Symmetry [?]. Synapse [?]. Syndrome [?]. Synthesis [?, ?, ?, ?]. Synthetic [?, ?]. System [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Systems [?, ?, ?, ?, ?].

Tablets [?]. Tadpole [?]. Takeoff [?]. Tar [?]. Targets [?]. Tartaria [?]. Taste [?]. Taxonomy [?]. Tchoga [?]. Teacher [?]. Teaching [?]. Tears [?]. Technological [?]. Technology [?, ?, ?, ?, ?, ?]. Tektites [?, ?, ?]. Telephone [?]. Telescope [?]. Temperature [?, ?]. Temperatures [?, ?]. Tension [?]. Teotihuacán [?]. Teredo [?]. Term [?]. Termite [?]. Terrestrial [?]. Territorial [?]. Test [?]. Tetanus [?]. Tetrodotoxin [?]. Texture [?]. Thalidomide [?]. Themistocles [?]. Theorems [?]. Theory [?, ?, ?, ?, ?, ?]. Thermal [?, ?]. Thermography [?]. Thermomagnetic [?]. Thermostat [?]. Things [?, ?]. Third [?, ?]. Third-Generation [?]. Three [?, ?, ?, ?]. Three-Dimensional [?, ?]. Three-Pigment [?]. Thumb [?]. Thymus [?, ?]. Thyroid [?]. Tides [?]. Tilapia [?]. Time [?, ?, ?, ?]. Time-Sharing [?]. Tom [?]. Tombs [?]. Tools [?, ?]. Topology [?]. Touch [?]. Toxic [?]. Toxin [?]. Trachoma [?]. Tracks [?]. Trade [?].

Trading [?]. Traffic [?, ?]. Transdetermination [?]. Transducers [?]. Transfer [?]. Transform [?]. Transformation [?, ?]. Translation [?, ?]. Transmission [?, ?]. Transpacific [?]. Transplanted [?]. Transplants [?]. Transport [?, ?]. Transportation [?, ?, ?]. Treasure [?]. Trees [?]. Triode [?]. Truth [?, ?]. Tube [?, ?]. Tumor [?, ?]. Tunnel [?]. Turbines [?]. Turkey [?, ?]. Turning [?]. Turtle [?]. Two [?, ?, ?]. Two-Mile [?]. Two-Neutrino [?]. Two-Phase [?]. Tycho [?]. Typesetting [?].

U.S. [?, ?, ?, ?, ?]. U.S.S.R. [?, ?]. Uganda [?]. Ultrahigh [?, ?]. Ultrahigh-Speed [?]. Ultrasonics [?]. Ultrasound [?]. Ultrastrong [?]. Ultraviolet [?, ?]. Undercooling [?]. Underground [?, ?]. Unorthodox [?]. Untouchables [?]. Upper [?, ?]. Urban [?, ?]. Urbanization [?]. Use [?]. Useful [?]. Uses [?, ?, ?, ?, ?].

Vacuum [?]. Valley [?]. Vaulting [?]. Vehicular [?]. Velocities [?]. Venous [?]. Venus [?]. Vertical [?]. Vertical-Takeoff [?]. Vessel [?]. Vibrating [?]. Vibration [?]. Vibrations [?]. Vikings [?]. Village [?, ?, ?]. Violations [?]. Violence [?]. Violins [?]. Virus [?, ?, ?, ?, ?, ?, ?]. Viruses [?, ?, ?, ?, ?, ?]. Vision [?, ?, ?, ?, ?]. Visual [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Vitro [?]. Volta [?]. Voltage [?]. Voodoo [?]. Vortex [?]. Voyage [?, ?].

Walking [?]. Wall [?]. War [?, ?]. was [?]. Wasps [?]. Water [?, ?, ?, ?, ?, ?, ?]. Water-Breathing [?]. Waves [?, ?, ?, ?, ?, ?]. Wear [?]. Weather [?, ?, ?, ?]. Webs [?]. Weddell [?]. West [?]. Whales [?]. Wheat [?]. Wheeled [?]. Whiskers [?]. Wild [?]. Wildlife [?]. William [?]. Wind [?]. Winds [?]. Wine [?]. Wires [?]. Withering [?]. without [?]. Women [?]. Wood [?]. work [?]. World [?, ?, ?, ?, ?, ?, ?]. Worm [?]. Wound [?].

X [?, ?, ?]. X-Ray [?, ?, ?].

Yachts [?]. Years [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Years [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Yeasts [?]. Yellowstone [?]. York [?]. Youngest [?].

Zanbil [?]. Zero [?]. Ziggurat [?]. Zodiacal [?]. Zone [?]. Zulu [?].